### v1

@@ Abstract.
Modern social media feeds use predictive models to maximize engagement, often misaligning how people consume content with how they wish to. We introduce Bonsai, a system that enables people to build personalized and intentional feeds. Bonsai implements a platform-agnostic framework comprising Planning, Sourcing, Curating, and Ranking modules. Altogether, this framework allows users to express their intent in natural language and exert fine-grained control over a procedurally transparent feed creation process. We evaluated the system with 15 Bluesky users in a two-phase, multi-week study. We find that participants successfully used our system to discover new content, filter out irrelevant or toxic posts, and disentangle engagement from intent, but curating intentional feeds required participants to exert more effort than they are used to. Simultaneously, users sought system transparency mechanisms to trust and effectively use intentional, personalized feeds. Overall, our work highlights intentional feedbuilding as a viable path beyond engagement-based optimization.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
At the heart of modern social media feeds are predictive models designed to maximize user engagement (Meta, 2025; Zhao et al., 2019; Liu et al., 2022; Cunningham et al., 2024). These algorithms analyze vast amounts of behavioral data to forecast which posts a user is most likely to view, like, share, or comment on (Narayanan, 2023). Platforms like Bluesky, X, Facebook, Instagram, and TikTok rely on this approach to rank and personalize content, effectively shaping each user’s feed into a unique, constantly adapting experience. This formula has clear benefits: by surfacing content that aligns with users’ revealed preferences, engagement-driven ranking can make apps feel more enjoyable and relevant (u/solutioneering, 2021). Past work has shown, for example, that users do indeed spend more time in feeds curated by these predictive systems over chronological ones (Guess et al., 2023).
@@
However, many users experience a disconnect between how they use social media in “the spur of the moment” and how they wish they used it in retrospect (Narayanan, 2023). In behavioral science lingo (Kahneman, 2011), this dichotomy reflects a tension between System 1 (fast, intuitive, and reactive) and System 2 (slower, deliberate, and reflective). While scrolling through a feed, users are often driven by immediate rewards like novelty or emotional stimulation. Afterwards, however, users may feel their usage did not align with deeper goals, such as staying informed or connecting meaningfully to others (Tran et al., 2019a; Kleinberg et al., 2024; Milli et al., 2025).
This gap between use and intent highlights a deeper challenge: users often lack the tools to align real-time experience with long-term goals (Sapacz et al., 2016; Marino et al., 2018; Alter, 2018). This gap may be a consequence of the traditional machine learning paradigm, requiring labeled training data, carefully engineered features, and task-specific models. Since social media platforms had vastly more data on implicit feedback than explicit feedback from users, this paradigm triumphed in practice. This idea is neatly described by Google employees describing YouTube’s recommender system: “Although explicit feedback mechanisms exist on YouTube (thumbs up/down, in-product surveys, etc.), we use the implicit feedback of watches to train the model, where a user completing a video is a positive example. This choice is based on the orders of magnitude more implicit user history available, allowing us to produce recommendations deep in the tail where explicit feedback is extremely sparse.” (Covington et al., 2016).
@@
In that context, language models (LMs) may change the rules of the game. Past work in HCI has explored various approaches to increase user agency—from reducing information overload (Bernstein et al., 2010) and creating friction to prevent mindless scrolling (Grüning et al., 2023; Hiniker et al., 2016), to helping users manually curate chronological feeds (Liu et al., 2025) or blocklists (Jhaver et al., 2018). However, these solutions often require significant user effort or sacrifice the benefits of algorithmic curation entirely. Unlike traditional machine learning systems or these manual interventions, LMs can use natural language instructions and adapt to different contexts without task-specific retraining, large labeled datasets, or constant user maintenance (Kojima et al., 2022; Wang et al., 2025). We hypothesize that these capabilities enable easy-to-use, intent-driven personalization. Instead of relying solely on engagement signals inferred from past behavior, recommender systems and feed algorithms can now incorporate direct user input about what they want to see. Further, these social media feeds allow more granular and transparent control, as prompts used by LMs directly specify what content gets sorted into the feed, and can be tweaked and adjusted.
@@
To investigate the feasibility and characteristics of intentional, personalized feeds, we introduce Bonsai, a system designed to increase user agency by helping them create personalized feeds that respond to intentions communicated in transparent, natural language prompts. We implemented Bonsai for Bluesky, an open and decentralized social media platform, as illustrated in Fig. 1, and conducted a two-phased interview study ($n=15$) to understand how users design and experience intentional feeds. Given that Bonsai explores a new design space, we evaluate it “in a holistic fashion on the basis of what [the system] makes possible,” focusing on “how compelling, how richly painted, and how informed is [the system’s] vision.” (Wobbrock, 2012).
@@
Bonsai was guided by three design goals: intent fidelity, ensuring that explicitly stated user goals govern feed generation; controllability, lowering the cost of preference articulation so users can flexibly adjust feeds without specialized expertise; and procedural transparency, making the process of sourcing, filtering, and ranking content visible and understandable. Together, these goals aim to shift personalization away from opaque engagement optimization toward feeds intentionally authored by users. To operationalize these goals, Bonsai implements a platform-agnostic four-stage. First, in the Planning stage, an LM agent parses the user’s natural language description into candidate sources and inclusion/exclusion criteria; Second, in the Sourcing stage, we fetch posts from feeds, lists, accounts, hashtags, or search queries from platform APIs, LMs, and previously scraped databases; Third, in the Curation stage, an LM agent scores candidate posts against user-stated preferences. Finally, in the Ranking stage, posts are ordered based on user-specified trade-offs among relevance, recency, and popularity. The result is not a static explanation of a black-box feed, but a transparent, editable configuration in which users iteratively co-author the algorithm that shapes their social media experience.
@@
Our findings suggest that intentional, personalized feeds help participants discover relevant content, focus on desired information, and avoid overload, ultimately strengthening their sense of agency. At the same time, these feeds demand greater effort to construct and heighten user expectations that the resulting content will faithfully reflect their stated intentions. While some users find excitement and fulfillment in building intentional social feeds, others struggle with the cognitive load associated with the task.
@@
Through this work, we aim to concretely demonstrate the feasibility of intentional, personalized social media through natural language. Overall, we contribute
@@
    Bonsai, a system that empowers intentional and personalized feedbuilding through explicit natural language input;
@@
    A content curation pipeline implemented for Bluesky, coordinating planning, sourcing, curating, and ranking through language models;
  3. (3)
    Results from a multi-week field study evaluating Bonsai’s potential to increase user agency and promote intentional social media use; and
@@ 2.1. Aligning Feeds with Explicit User Goals
Yet maximizing engagement often conflicts with users’ explicit goals. Experimental and observational work shows that engagement-driven feeds can promote addictive and regrettable use (Tran et al., 2019a; Cho et al., 2021), exacerbate compulsive habits (Tran et al., 2019b), and surface uncivil or emotionally charged content that users later report not wanting to see (Guess et al., 2023; Milli et al., 2025). Deactivation studies find that participants report improved well-being after leaving social media altogether (Tromholt, 2016; Allcott et al., 2020, 2025), which may not be all that surprising since previous work suggests that around one in four people worldwide exhibit signs of problematic or addictive use (Cheng et al., 2021). These patterns have fueled broad concerns about the societal harms of engagement-optimized feeds (Kleinberg et al., 2024; Harris, 2021).
@@ 2.2. User Agency and Control in Feeds
Early interventions primarily focused on when and how much technology was used. For example, MyTime provided lightweight scheduling to reduce time spent in “poor-use” apps while preserving “good-use” ones (Hiniker et al., 2016), NUGU leveraged group accountability to curb overuse (Ko et al., 2015), and OneSec inserted intentional delays to disrupt habitual app opening (Grüning et al., 2023). These systems were effective at mitigating compulsive and ‘undesired’ use, for instance MyTime reduced usage on apps that participants felt were a poor use of time by 21% while keeping the usage of ‘useful’ apps unchanged. Yet, note that these systems intervene at the temporal level of use, rather than shaping the content that fills feeds.
@@
A parallel line of work has explored richer, user-driven controls within social media feeds themselves. Eddi enabled interactive topic browsing by clustering feed items into trending topics; users found browsing to be more efficient, enjoyable, and less overwhelming than chronological interfaces, often noting they could “find things […] faster” and avoid irrelevant content (Bernstein et al., 2010); Cura, on the other hand, uses transformer-based modeling to predict which posts a designated curator would endorse based on community upvotes; evaluations demonstrated that its curated feeds accurately reflected curator preferences, adapting noticeably when curators changed, and halving antisocial behavior without extra moderation effort (He et al., 2023). Locus scaffolds intentional entry points such as choosing a session purpose (e.g., ”learning” vs. ”socializing”), particularly helping younger users select content aligned with their intentions (Davis et al., 2023). These tools demonstrate how systems can extend beyond managing time to empower users with nuanced control over what they see. Last, besides individual systems, studies of decentralized platforms show how users manually combine follows, lists, and custom rules to assemble personally meaningful streams (Liu et al., 2025).
@@
With the increasing capabilities of language models, a growing body of work has explored natural language input as a mechanism for communicating preferences to algorithms (Zhou et al., 2024; Gao et al., 2021). Wang et al. (2025) compared three strategies for end-user content curation: example labeling, rule writing, and prompting language models. They found that natural language prompts enabled users to rapidly bootstrap personalized classifiers with strong performance. This insight aligns with a broader tradition of conversational recommendation, where preference elicitation through naturalistic questions often outperforms rigid item-attribute queries (Kostric et al., 2021) and recent work showing that language models can be integrated directly into recommendation pipelines (Sanner et al., 2023).
@@
This prior work motivates Bonsai to shift from regulating when people use technology to shaping what they see. Unlike previous work, Bonsai elicits user goals directly through natural language and makes each stage of feed construction explicit and editable. In doing so, it operationalizes user agency not just by letting people tune or explore feeds, but by enabling them to design the very algorithms that govern their online experience.
@@ 2.3. Transparency and Algorithmic Feeds
Traditionally, transparency in feed algorithms has focused on helping users make sense of a black-box model after the fact. For example, surfacing factors correlated with recommendations or disclosing which signals influence ranking (Tintarev and Masthoff, 2015). Other efforts have emphasized exposing the assumptions and data inputs used to personalize feeds (rader2015beliefs; Rader et al., 2018). In contrast, Bonsai offers a distinct approach: rather than retrofitting explanations onto engagement-based algorithms, it makes the feed construction pipeline itself transparent. This shifts the focus away from what Ananny and Crawford (2018) critique as “black-box” ideals of transparency  toward a model in which users gain clarity by actively authoring the algorithmic process that shapes their feeds.
Figure 2. Positioning of different social media feeds along two dimensions: intentional vs. emergent and generic vs. personalized. Twitter’s old chronological feed is generic and intentional; Reddit’s “Hot” feed is generic and emergent, TikTok’s “For You” feed is personalized and emergent; finally, Bonsai represents an intentional and personalized approach.
[…]
@@
Bonsai explores the design space for social media feeds to be both intentional and personalized. Before detailing the specific design goals, we distinguish social media feeds alongside two dimensions: intentional vs. emergent and generic vs. personalized.
Emergent vs. Intentional Feeds. A feed is intentional when people’s experience is shaped by the goals they explicitly articulate, for example by selecting topics or accounts to feature. Conversely, emergent feeds derive from patterns of everyday use, with the system inferring intent from engagement and other in-platform behaviors.
@@
We illustrate this taxonomy with three prominent social media feeds in Fig. 2. TikTok’s “For You Feed” is emergent and personalized; recommendations are primarily driven by predicted user engagement, and your engagement with content strongly shapes the content you are shown (Vombatkere et al., 2024). Reddit’s “Hot Feed” is emergent and generic; the feed is created with a simple formula that considers recency, but also engagement from others, including comment rate and upvote score (u/solutioneering, 2021). Old Twitter’s “Chronological Feed” is explicit and generic; the feed draws from sources that users explicitly choose to follow, and rank their posts based on general criteria (recency) that do not consider engagement signals (DeVito et al., 2017). By contrast, Bonsai occupies the explicit and personalized quadrant, enabling users to author feeds that are tailored to their stated intentions while still adapting to their individual preferences.

### v2

@@ Abstract.
Social media feeds use predictive models to maximize engagement, often misaligning how people consume content with how they wish to. We introduce Bonsai, a system that enables people to build personalized and intentional feeds. Bonsai implements a platform-agnostic framework comprising Planning, Sourcing, Curating, and Ranking modules. This framework allows users to express their intent in natural language and exert fine-grained control over a procedurally transparent feed creation process. We evaluated the system with 15 Bluesky users in a two-phase, multi-week study. We find that participants successfully used our system to discover new content, filter out irrelevant or toxic posts, and disentangle engagement from intent, but curating intentional feeds required more effort than they are used to. Simultaneously, users sought system transparency mechanisms to effectively use (and trust) intentional, personalized feeds. Overall, our work highlights intentional feedbuilding as a viable path beyond engagement-based optimization.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791855isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in HCI
Work published at CHI 2026. Please cite accordingly.
@@ 1. Introduction
At the heart of modern social media feeds are predictive models designed to maximize user engagement (Meta, 2025; Zhao et al., 2019; Liu et al., 2022). These algorithms analyze vast amounts of behavioral data to forecast which posts a user is most likely to view, like, share, or comment on (Narayanan, 2023). Platforms like Bluesky, X, Facebook, Instagram, and TikTok rely on this approach to rank and personalize content, effectively shaping each user’s feed into a unique, constantly adapting experience. This formula aligns tightly with advertising-based business models that monetize attention and impressions, giving platforms strong incentives to optimize for engagement (Kim, 2017). It also has clear short-term benefits: engagement-driven ranking can make apps feel enjoyable and relevant (u/solutioneering, 2021), and users spend more time in feeds curated by these predictive systems than in chronological ones (Guess et al., 2023).
@@
At the same time, many users experience a disconnect between how they use social media in “the spur of the moment” and how they wish they used it in retrospect (Narayanan, 2023). In behavioral science lingo (Kahneman, 2011), this dichotomy reflects a tension between System 1 (fast, intuitive, and reactive) and System 2 (slower, deliberate, and reflective). While scrolling through a feed, users are often driven by immediate rewards like novelty or emotional stimulation. Afterwards, however, users may feel their usage did not align with deeper goals, such as staying informed or connecting meaningfully to others (Tran et al., 2019b; Kleinberg et al., 2024; Milli et al., 2025). This gap between use and intent contributes to compulsive habits (Tran et al., 2019a), problematic or addictive use (Marino et al., 2018; Sapacz et al., 2016), and even decisions to deactivate platforms altogether (Allcott et al., 2020; Tromholt, 2016).
But how can users increase their agency while using social media? Past work in HCI and CSCW has explored a range of approaches to help people feel more in control of their feeds and their time: from topic-based interfaces that reduce information overload (Bernstein et al., 2010) to introducing friction to disrupt mindless scrolling (Hiniker et al., 2016; Grüning et al., 2023). On decentralized social media platforms in particular, communities have built customization tools, such as user-authored feeds on Bluesky (e.g., SkyFeed), that let people specify rules, keywords, and filters to shape what appears in their timelines. These systems show that some users are motivated to take a more active role in configuring their feeds, and that even relatively simple customization can make platforms feel more aligned with their goals.
Yet existing approaches face important limitations. Time management and friction tools can reduce overuse, but they are blunt instruments, preventing use altogether or restricting access to entire apps, rather than helping users transform how they use those apps. Rule-based customization tools can support fine-grained control, but, in practice, require substantial effort and technical understanding. Users must translate high-level goals (e.g., “see more thoughtful posts from climate policy experts”) into low-level rules, lists, and filters, and must continually adjust these configurations as their interests and the platform evolve. As a result, many customization tools remain niche, and engagement-optimized feeds continue to dominate the default user experience, powered by the vast amounts of implicit feedback signals that platforms have (Covington et al., 2016).
@@
In that context, language models (LMs) may change the rules of the game. Unlike traditional machine learning systems or manual interventions, LMs can follow natural language instructions and adapt to different contexts without task-specific retraining, large labeled datasets, or constant user maintenance (Kojima et al., 2022; Wang et al., 2025). These capabilities may enable easy-to-use, intent-driven personalization. Instead of relying on engagement signals inferred from past behavior, recommender systems and feed algorithms can now incorporate direct user input about what they want to see. They could do so at a level of specificity (e.g., tone, sources, trade-offs between relevance and serendipity) that would be cumbersome to encode as rules. Further, such feeds can offer more granular, legible control: prompts and configuration settings directly specify which content enters the feed and can be inspected, discussed, and iteratively refined.
@@
To investigate the feasibility and characteristics of intentional, personalized feeds, we adopt an exploratory, systems-based approach. Broadly, we ask: How do LM-mediated feed builders reshape how people author, experience, and sustain personalized social media feeds? We investigate this overarching direction with Bonsai, a system designed to increase user agency by helping them create personalized feeds that respond to intentions communicated in transparent, natural language prompts. We implemented Bonsai for Bluesky, an open and decentralized social media platform, as illustrated in Fig. 1, and conducted a two-phased interview study ($n=15$) to understand how users design and experience intentional feeds. Given that Bonsai explores a new design space, we evaluate it “in a holistic fashion on the basis of what [the system] makes possible,” focusing on “how compelling, how richly painted, and how informed is [the system’s] vision.” (Wobbrock, 2012).
@@
Bonsai was guided by three design goals: intent fidelity, ensuring that explicitly stated user goals govern feed generation; ease of use, lowering the cost of preference articulation so users can flexibly adjust feeds without specialized expertise; and procedural transparency, making the sourcing, filtering, and ranking of content visible and understandable. Together, these goals aim to shift personalization away from engagement optimization toward feeds intentionally authored by users. To operationalize these goals, Bonsai implements a platform-agnostic four-stage framework. First, in the Planning stage, an LM agent parses the user’s natural language description into candidate sources and inclusion/exclusion criteria; Second, in the Sourcing stage, we fetch posts from feeds, lists, accounts, hashtags, or search queries from platform APIs, LMs, and previously scraped databases; Third, in the Curation stage, an LM agent scores candidate posts against user-stated preferences. Finally, in the Ranking stage, posts are ordered based on user-specified trade-offs among relevance, recency, and popularity. The result is an editable configuration in which users iteratively co-author the algorithm that shapes their social media experience.
@@
Our findings suggest that intentional, personalized feeds help participants discover relevant content, focus on desired information, and avoid overload, ultimately strengthening their sense of agency. The LM-mediated nature of Bonsai enabled participants to express nuanced preferences, convey social media intentions through precise post specifications, and allowed novice users to onboard easily to custom feeds. At the same time, these feeds demand greater effort to construct and heighten user expectations that the resulting content will faithfully reflect their goals; while some users find excitement and fulfillment in building intentional social feeds, others struggle with the cognitive load associated with the task. Participant feedback highlights the promise of hybrid approaches that combine intentional and engagement-based feeds.
@@
Through this work, we aim to concretely demonstrate the feasibility of intentional, personalized social media through natural language. Overall, we contribute:
@@
    Bonsai, a system that empowers intentional and personalized feedbuilding through explicit natural language input. To the best of our knowledge, Bonsai is the first end-user-facing system using language models to enable people to author and control social media feeds in natural language.
@@
    Results from a multi-week field study with 15 Bluesky users that characterize how people actually author, experience, and sustain LM-mediated intentional feeds with Bonsai
@@ 2.1. Aligning Feeds with Explicit User Goals
Yet maximizing engagement often conflicts with users’ explicit goals. Experimental and observational work shows that engagement-driven feeds can promote addictive and regrettable use (Tran et al., 2019b; Cho et al., 2021), exacerbate compulsive habits (Tran et al., 2019a), and surface uncivil or emotionally charged content that users later report not wanting to see (Guess et al., 2023; Milli et al., 2025). Deactivation studies find that participants report improved well-being after leaving social media altogether (Tromholt, 2016; Allcott et al., 2020, 2025), which may not be all that surprising since previous work suggests that around one in four people worldwide exhibit signs of problematic or addictive use (Cheng et al., 2021). These patterns have fueled broad concerns about the societal harms of engagement-optimized feeds (Kleinberg et al., 2024; Harris, 2021).
@@
In parallel, a growing body of work in Human–AI Interaction shows that users do not simply consume algorithmic outputs, but actively try to steer and collaborate with AI systems. Interactive and “teachable” systems let people correct or refine model behavior through examples, labels, or natural-language feedback, and find that users are most willing to provide such input when it clearly influences outcomes and remains lightweight (Amershi et al., 2014; Ramos et al., 2020; Feng et al., 2024). In this view, aligning feeds with explicit goals is part of a broader shift toward mixed-initiative systems, where humans articulate high-level intentions while the AI infers low-level details and proposes refinements over time (Horvitz, 1999).
@@
Table 1. Description of systems that support more intentional technology use and feed control.
System Mechanism Level of abstraction for users
MyTime (Hiniker et al., 2016) User-defined schedules to reduce time spent in “poor-use” apps while preserving “good-use” ones. App-level time blocks.
NUGU (Ko et al., 2015) Group-based accountability and shared goals to curb the overuse of smartphones. High-level usage goals and social check-ins. App-level time blocks.
OneSec (Grüning et al., 2023) Inserts intentional delays and breathing prompts before opening selected apps to interrupt habitual use. Simple app-level switches.
Eddi (Bernstein et al., 2010) Clusters feed items into topical groups and lets users browse and filter by topic. Topic-level browsing and filtering; lightweight manual exploration.
Cura (He et al., 2023) Predicts which posts a designated curator would endorse from community upvotes and generates curated feeds. Community-level preferences inferred by the model.
Locus (Davis et al., 2023) Prompts users to choose a session purpose (e.g., learning vs. socializing), and reflect upon usage High-level session goals and reflections.
Skyfeed and Graze11footnotemark: 1 Visual builders and filters allow users to define custom feed rules over follows, keywords, and interactions. Low-level rules, filters, and regex.
Bonsai High-level goals described in natural language. Language models map these goals to editable, language-first configurations for sourcing, curating, and ranking content. High-level natural-language prompts.
The table compares eight systems that support more intentional technology use and feed control along three columns: System, Mechanism, and Level of abstraction for users. It lists MyTime, NUGU, OneSec, Eddi, Cura, Locus, Skyfeed & Graze, and Bonsai. Earlier systems focus on scheduling or friction (e.g., MyTime, OneSec), group accountability (NUGU), topic or curator-based feeds (Eddi, Cura), or session goals (Locus), with user interaction at the app, topic, or rule level. Skyfeed and Graze require users to specify low-level rules and filters, whereas Bonsai is distinguished as using high-level natural-language prompts that are mapped by language models into configurations for sourcing, curating, and ranking content.
@@ 2.2. User Agency and Control in Feeds
Early interventions primarily focused on when and how much technology was used. For example, MyTime provided lightweight scheduling to reduce time spent in “poor-use” apps while preserving “good-use” ones (Hiniker et al., 2016), NUGU leveraged group accountability to curb overuse (Ko et al., 2015), and OneSec inserted intentional delays to disrupt habitual app opening (Grüning et al., 2023). These systems were effective at mitigating compulsive and ‘undesired’ use. For instance MyTime reduced usage on apps that participants felt were a poor use of time by 21% while keeping the usage of ‘useful’ apps unchanged. Yet, note that these systems intervene at the temporal level of use, rather than shaping the content that fills feeds.
@@
A parallel line of work has explored richer, user-driven controls within social media feeds themselves. Eddi enabled interactive topic browsing by clustering feed items into trending topics; users found browsing to be more efficient, enjoyable, and less overwhelming than chronological interfaces, often noting they could “find things […] faster” and avoid irrelevant content (Bernstein et al., 2010); Cura, on the other hand, uses transformer-based modeling to predict which posts a designated curator would endorse based on community upvotes; evaluations demonstrated that its curated feeds accurately reflected curator preferences, adapting noticeably when curators changed, and halving antisocial behavior without extra moderation effort (He et al., 2023). Locus scaffolds intentional entry points such as choosing a session purpose (e.g., ”learning” vs. ”socializing”), particularly helping younger users select content aligned with their intentions (Davis et al., 2023). On Bluesky, we note two third-party tools: Skyfeed and Graze. These tools let users author custom feed logic through visual rule builders, regular expressions, and composable filters, [^1] E.g., https://skyfeed.app, https://www.graze.social. exemplifying the promise of “algorithmic choice,” while still requiring users to think in terms of low-level rules.
@@
With the increasing capabilities of language models, a growing body of work has explored natural language input as a mechanism for communicating preferences to algorithms (Zhou et al., 2024; Gao et al., 2021). Wang et al. (2025) compared three strategies for end-user content curation: example labeling, rule writing, and prompting language models. They found that natural language prompts enabled users to rapidly bootstrap personalized classifiers with strong performance. This insight aligns with a broader tradition of conversational recommendation, where preference elicitation through naturalistic questions often outperforms rigid item-attribute queries (Kostric et al., 2021) and recent work showing that language models can be integrated directly into recommendation pipelines (Sanner et al., 2023). However, despite this technical promise, LLM-powered recommenders have not yet seen widespread deployment in production feeds, in part due to cost, latency, and safety challenges at scale. Major platforms are only beginning to experiment with these ideas in their core products (Perez, 2025; Silberling, 2025; Musk, 2025)
@@
This prior work, summarized in Table 1, motivates Bonsai to shift from regulating when people use technology to shaping what they see. Unlike prior systems that either operate at the temporal level (e.g., scheduling and friction tools), or require users to hand-craft low-level rules and filters, Bonsai elicits user goals directly through natural language and makes each stage of feed construction explicit and editable.
@@ 2.3. Transparency and Algorithmic Feeds
A parallel, broader line of work in Human–AI Interaction examines how transparency shapes ongoing collaboration between people and AI systems. Rather than treating explanations as purely informational, this research shows that users appropriate them to calibrate trust, adjust strategies, and figure out how to “help” the system behave as desired (Kim et al., 2023; Liao and Vaughan, 2023). For example, in an image-recognition setting, Kim et al. (2023) find that participants use model explanations to improve their own inputs and to diagnose when the AI has misunderstood a case, not just to verify predictions. More broadly, human-centered explainability work argues that transparency is most beneficial when it is tightly coupled to concrete levers for action, so that users can translate understanding into effective interventions (Ehsan and Riedl, 2020; Liao and Varshney, 2021).
Transparency in feed algorithms has focused on helping users make sense of a black-box model after the fact, e.g., surfacing factors correlated with recommendations or disclosing which signals influence ranking (Tintarev and Masthoff, 2015). Other efforts have emphasized exposing the assumptions and data inputs used to personalize feeds (Rader and Gray, 2015; Rader et al., 2018). In contrast, Bonsai offers a distinct approach: rather than retrofitting explanations onto engagement-based algorithms, it makes the feed construction pipeline itself transparent. This shifts the focus away from what Ananny and Crawford (2018) critique as “black-box” ideals of transparency  toward a model in which users gain clarity by actively authoring the algorithmic process that shapes their feeds.
Figure 2. Positioning of different social media feeds along two dimensions: intentional vs. emergent and generic vs. personalized. These axes are introduced in this paper as an analytic framing and do not reflect platform categories. Twitter’s old chronological feed is generic and intentional; Reddit’s “Hot” feed is generic and emergent; TikTok’s “For You” feed is personalized and emergent. Bonsai appears in the remaining quadrant as a system for constructing feeds, not a platform.
@@
Bonsai explores the design space for social media feeds to be both intentional and personalized. Before detailing the specific design goals, we distinguish social media feeds according to two dimensions: intentional vs. emergent and generic vs. personalized.
Intentional vs. Emergent Feeds. A feed is intentional when people’s experience is shaped by the goals they explicitly articulate, e.g., by selecting topics or accounts to feature. Conversely, emergent feeds derive from patterns of everyday use, with the system inferring intent from engagement and other in-platform behaviors.
@@
We illustrate this taxonomy with three prominent social media feeds in Fig. 2. TikTok’s “For You Feed” is emergent and personalized; recommendations are primarily driven by predicted user engagement, and your engagement with content strongly shapes the content you are shown (Vombatkere et al., 2024). Reddit’s “Hot Feed” is emergent and generic; the feed is created with a simple formula that considers recency, but also engagement from others, including comment rate and upvote score (u/solutioneering, 2021). Old Twitter’s “Chronological Feed” is intentional and generic; the feed draws from sources that users explicitly choose to follow, and ranks their posts based on general criteria (recency) that do not consider engagement signals (DeVito et al., 2017). In contrast, Bonsai is in the intentional and personalized quadrant, enabling feeds tailored to stated intentions while still adapting to individual preferences.